OpenStudy (chris215):

If f(-1) = -3 and f'(x)=(4x^2)/(x^3+3), which of the following is the best approximation for f(-1.1) using local linearization?

OpenStudy (chris215):

-7.2 2.8 -1.2 -3.2

OpenStudy (jack1):

hey, just curious... do u understand the theory of solving ODE's (1st and 2nd)?

OpenStudy (jack1):

not preaching, just if u do, i can help, if not.... it's just after 5:30am here, and i've been awake all night, so imma go to sleep pretty soon

OpenStudy (chris215):

ohh ok I dont but thats ok

OpenStudy (jack1):

ok, im going to assume that's a yes? 1st order linear ODE exact solution is: f(x) = 1/3 (4 log(x^3+3)-9-(4 log(2))) so plug x = -1.1 into this solve. u ok with that?

OpenStudy (jack1):

sorry... @zepdrix are you able to assist @chris215 with the step by step, I'm pretty much crashing hard right now

zepdrix (zepdrix):

So for linear approximation we use something like this, ya?\[\large\rm f(x)\approx f(x_0)+f'(x)(x-x_o)\] To approximate our function f at some x, we calculate a nearby point f(x_o) and then add some stuff to it. For this problem, our point that we're approximating is at \(\large\rm x=-1.1\). We're using a place nearby to approximate \(\large\rm x_o=-1\). So for our approximation, f(-1.1) is going to be about this much,\[\large\rm f(-1.1)\approx f(-1)+f'(-1.1)(-1.1-(-1))\]

zepdrix (zepdrix):

We can simplify the last part in the brackets,\[\large\rm f(-1.1)\approx f(-1)-0.1\cdot f'(-1.1)\]

zepdrix (zepdrix):

They gave us information to figure the rest out, f(-1)=-3 and f'(x)=4x^2/(x^3+3) So plug in your -3 in place of f(-1), and also, evaluate f'(x) at x=-1.1, and plug it in.

OpenStudy (zarkon):

\[\Large L(x)=f(a)+f'(a)(x-a)\] \[\Large L(-1.1)=f(-1)+f'(-1)(-1.1-(-1))\]

OpenStudy (zarkon):

notice that \(f'(x)\) is evaluated at -1 and not -1.1

zepdrix (zepdrix):

woops :))

OpenStudy (chris215):

Thanks guys :))